OpenStudy (gabylovesyou):

Where are the x-intercepts for f(x) = 4 cos(2x − π) from x = 0 to x = 2π? x = pi over 4, x = 3 pi over 4, x = 5 pi over 4 x = pi over 2, x = 3 pi over 2 x = pi over 4, x = 3 pi over 4, x = 5 pi over 4, x = 7 pi over 4 x = pi over 2, x = 3 pi over 2, x = 5 pi over 2

OpenStudy (gabylovesyou):

@agent0smith

OpenStudy (agent0smith):

x-int means y=0, so you have to solve 4 cos(2x − π) = 0

OpenStudy (gabylovesyou):

x = 3pi/ 4

OpenStudy (agent0smith):

There's gonna be more than one.

OpenStudy (agent0smith):

Think about where cos theta = 0. You'll have to list several solutions, or use the +2*k*pi, where k is an integer. Then set 2x − π equal to those, and solve for x.

OpenStudy (gabylovesyou):

is the one i said one of the options atleast ? so i know it could be A or C ..

OpenStudy (agent0smith):

Yeah the one you said is right

OpenStudy (gabylovesyou):

ok i dont know..

OpenStudy (gabylovesyou):

I'm going to take a guess and say A.

OpenStudy (agent0smith):

You could just plug in the other options and see if they work too... or use the fact cos theta equals zero when theta is either pi/2 + 2 k pi or 3pi/2 + 2 k pi then 2x − π = pi/2 +2 k pi or 2x − π = 3pi/2 +2 k pi and solve for x, and plug in values of k like 0, 1, 2 ,3... until you have all the ones from 0 to 2pi

OpenStudy (gabylovesyou):

Ohhhh... ok thank you.